An automobile manufacturer claims that its jeep has a 35.6 miles/gallon (MPG) rating. An independent testing firm has been contracted to test the MPG for this jeep since it is believed that the jeep has an incorrect manufacturer's MPG rating. After testing 240 jeeps, they found a mean MPG of 35.2. Assume the variance is known to be 5.76. A level of significance of 0.05 will be used. Find the value of the test statistic. Round your answer to 2 decimal places. Enter the value of the test statistic.
Answer #1

Solution :

variance = 5.76

sigma = 2.4

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (35.2 - 35.6) / 2.4 / \sqrt 240

= -2.58

Test statistic = -2.58
Answer #2

An automobile manufacturer has given its jeep a 31.5 miles/gallon (MPG) rating. An independent testing firm has been contracted to test the actual MPG for this jeep since it is believed that the jeep has an incorrect manufacturer's MPG rating. After testing 210 jeeps, they found a mean MPG of 31.3. Assume the population variance is known to be 7.29. Is there sufficient evidence at the 0.05 level to support the testing firm's claim?

Step 2 of 6: 

Find the value of the test statistic. Round your answer to two decimal places.